Grade 8Mechanics

  • A simple pendulum has a time period T1 when on the earths surface , and T2 when taken to a height R above the earths surface where R is the radius of the earth . The value of T2/T1 is

  • 1

  • √2

  • 4

  • 2

mgayathri
9 Years agoGrade 8

2 Answers

DR STRANGE
9 Years ago
T=2pi*sqrt(l/g)
g is inversly proportional to (R+h)2 …..........h is height above the earth’s  surface
therefore T is proportional to R+h
T2/T1= (R+h)/R =2............ putting h=R
Adsxsc
8 Years ago
the time period of a pendulum on earth surface is T1 its time period at a height equal to radius of the earth is T2 then ratio of T1:T2 is